### 13 Mar 2026 v1

# Filtered Spectral Projection for Quantum Principal Component Analysis

Abstract
Quantum principal component analysis (qPCA) is commonly formulated as the extraction of eigenvalues and eigenvectors of a covariance-encoded density operator. Yet in many qPCA settings, the practical objective is simpler: projecting data onto the dominant spectral subspace. In this work, we introduce a projection-first framework, the Filtered Spectral Projection Algorithm (FSPA), which bypasses explicit eigenvalue estimation while preserving the essential spectral structure. FSPA amplifies any nonzero warm-start overlap with the leading principal subspace and remains robust in small-gap and near-degenerate regimes without inducing artificial symmetry breaking in the absence of bias. To connect this approach to classical datasets, we show that for amplitude-encoded centered data, the ensemble density matrix $\rho=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$ coincides with the covariance matrix. For uncentered data, $\rho$ corresponds to PCA without centering, and we derive eigenvalue interlacing bounds quantifying the deviation from standard PCA. We further show that ensembles of quantum states admit an equivalent centered covariance interpretation. Numerical demonstrations on benchmark datasets, including Breast Cancer Wisconsin and handwritten Digits, show that downstream performance remains stable whenever projection quality is preserved. These results suggest that, in a broad class of qPCA settings, spectral projection is the essential primitive, and explicit eigenvalue estimation is often unnecessary.
Quantum ML, Principal Component Analysis, Projection
preprint: AIP/123-QED
Quantum principal component analysis (qPCA) is one of the foundational ideas in quantum machine learning (QML) because it connects quantum linear-algebra methods to dimensionality reduction, feature extraction, and covariance-structure analysis in high-dimensional data.[1, 2, 3, 4, 5] Its importance is both conceptual and algorithmic: qPCA helped establish the view that quantum states can be treated as computational objects for learning tasks, rather than merely measured as passive outputs.[1]
Spectral primitives are central to quantum algorithms for learning, tomography, and simulation, with qPCA as a canonical example. The original Lloyd–Mohseni–Rebentrost construction shows that copies of a density matrix can be used to implement $e^{-i\rho t}$ and, via phase estimation, extract principal eigendirections in quantum form.[1] This viewpoint remains foundational because it reframes a quantum state as an active resource in its own analysis.
At the same time, two limitations have become increasingly clear. First, estimation-first pipelines inherit precision overhead from phase resolution and can become fragile when useful eigenvalues are uniformly compressed in magnitude, even at fixed ordering.[1, 6] Second, complexity claims based on idealized state-preparation access must be interpreted carefully: dequantization results show that part of the apparent exponential advantage can come from the access model rather than from a fundamentally different objective.[7] Together, these points motivate a sharper formulation of the task before choosing the spectral primitive.
Our starting observation is that many qPCA-motivated applications are projection tasks, not eigenvalue-estimation tasks. In near-degenerate regimes, individual leading eigenvectors are unstable and basis-dependent, whereas the dominant invariant subspace is stable and operationally meaningful. A projection-targeted method should therefore prioritize overlap amplification onto that subspace and avoid explicit dependence on global eigenvalue magnitude.
We introduce the Filtered Spectral Projection Algorithm (FSPA), an adaptive filter-and-renormalize iteration that amplifies dominant spectral support without explicitly encoding eigenvalues. At the state-update level, FSPA is equivalent to normalized power iteration with an adaptive schedule; we therefore do not claim improved gap-scaling asymptotics over that baseline.[8] The contribution is instead task-level: a projection-first quantum spectral primitive with built-in magnitude invariance and explicit handling of degenerate dominant spectra, naturally interpretable in block-encoding/QSVT language.[6, 9]
Main results are as follows. (i) FSPA is invariant under uniform spectral rescaling. (ii) Convergence remains gap-limited and follows the expected power-iteration-type complexity. (iii) Under degeneracy, convergence is to the dominant invariant subspace; with nonzero positive semidefinite symmetry breaking inside that subspace, the selected dominant direction is recovered. Numerical experiments reproduce the magnitude-collapse regime reported for estimation-first qPCA and show that FSPA removes this collapse while preserving the expected gap dependence.
Let
$\rho=\sum_{j=1}^{d}\lambda_{j}|\psi_{j}\rangle\langle\psi_{j}|$
[…]
  • Eigenvalue estimation: approximate $\lambda_{j}$ to prescribed precision.
FSPA targets dominant-subspace recovery and overlap amplification directly rather than eigenvalue estimation. The iteration used in this work is:
Algorithm 1 Filtered Spectral Projection Algorithm (FSPA)
[…]
2:State with amplified overlap on dominant eigenspace
[…]
5:Set amplification parameter $\beta\leftarrow 1$
6:for $t=1$ to $T$ do
7:  for $k=1$ to $\beta$ do
[…]
10:  end for
11:  $\beta\leftarrow 2\beta$
12:end for
[…]
The adaptive growth of $\beta$ yields progressively stronger filtering while the renormalization step removes global eigenvalue-scale dependence.[8]
At the state-update level, FSPA is mathematically equivalent to normalized power iteration applied with an adaptive schedule. This relation is intentional, not a hidden limitation. What is standard is the gap-limited amplification mechanism; what is new here is the projection-first formulation in qPCA settings where magnitude collapse of estimation-first methods is relevant.[6]
From a polynomial viewpoint, each round applies a low-degree spectral filter to $\rho$, and the adaptive schedule generates a sequence of increasing effective degrees. This interpretation aligns with block-encoding/QSVT realizations of polynomial operator transforms.[9, 10]
[…]
Figure 1: Schematic illustration of FSPA. Repeated application of $\rho$ amplifies components proportionally to $\lambda_{j}^{k}$, exponentially suppressing subdominant eigenspaces. Renormalization removes magnitude dependence and yields projection onto the dominant eigenspace.
The normalized iteration underlying FSPA has several structural properties that can be stated compactly.
Proposition 1 (Eigenvalue Magnitude Invariance)
Let $\rho$ be a Hermitian operator and let $c>0$. The normalized iterates produced by FSPA applied to $\rho$ and $c\rho$ are identical at every step. In particular, FSPA is invariant under uniform rescaling of the spectrum.
This first property is immediate from renormalization but central to the interpretation of FSPA as a projection-first primitive.
Proposition 2 (Gap-Dependent Bias Amplification)
Let $\rho$ be Hermitian with ordered eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\cdots$. If $\lambda_{1}>\lambda_{2}$ and the initial state has nonzero overlap with the dominant eigenvector, then the overlap produced by FSPA increases monotonically with the number of rounds. The convergence rate is governed by $\lambda_{1}/\lambda_{2}$, equivalently by the spectral gap. If $\lambda_{1}=\lambda_{2}$, convergence is only to the corresponding invariant subspace.
Proposition 3 (Subspace Convergence)
Let the dominant eigenspace of $\rho$ have degeneracy $R\geq 1$, i.e., $\lambda_{1}=\cdots=\lambda_{R}>\lambda_{R+1}$. Then, in the limit of infinite amplification rounds $T\to\infty$, FSPA produces a normalized state whose fidelity with the dominant invariant subspace approaches unity. Convergence to a unique eigenvector is not guaranteed unless degeneracy is lifted.
These propositions clarify the roles of scale invariance, spectral ratios, and degeneracy. The main quantitative guarantee is the following.
Theorem 1 (Oracle Complexity of FSPA)
Let $\rho$ be a Hermitian operator with spectral decomposition
$|,\rho=\sum_{j=1}^{d}\lambda_{j}\ket{\psi_{j}}\bra{\psi_{j}}$,
where $\lambda_{1}>\lambda_{2}\geq\cdots\geq 0$. Let the initial state be
[…]
with $a_{1}\neq 0$. Define the spectral ratio
$r:=\frac{\lambda_{2}}{\lambda_{1}}<1.$
[…]
$|\braket{\psi_{1}}{\phi_{k}}|^{2}\geq 1-\epsilon$
after at most
$\mathcal{O}\!\left(\frac{\log\!\left(\frac{1}{\epsilon}\right)+\log\!\left(\frac{1}{|a_{1}|^{2}}\right)}{\log(\lambda_{1}/\lambda_{2})}\right)$
applications of $\rho$.
Proof sketch: We expand the initial state in the eigenbasis,
[…]
and apply $\rho^{k}$:
[…]
Factoring out $\lambda_{1}^{k}$ gives
[…]
where $r_{j}=\lambda_{j}/\lambda_{1}\leq r<1$.
After normalization,
$F_{k}=\frac{|a_{1}|^{2}}{|a_{1}|^{2}+\sum_{j\geq 2}|a_{j}|^{2}r_{j}^{2k}}.$
Using $r_{j}\leq r$ gives
$F_{k}\geq\frac{|a_{1}|^{2}}{|a_{1}|^{2}+(1-|a_{1}|^{2})r^{2k}}.$
Imposing $F_{k}\geq 1-\epsilon$ yields
$r^{2k}\leq\frac{\epsilon}{1-\epsilon}\frac{|a_{1}|^{2}}{1-|a_{1}|^{2}}.$
Taking logarithms,
$k\geq\frac{\log\!\left(\frac{1-\epsilon}{\epsilon}\frac{1-|a_{1}|^{2}}{|a_{1}|^{2}}\right)}{2\log(\lambda_{1}/\lambda_{2})}.$
Therefore,
$\mathcal{O}\!\left(\frac{\log(1/\epsilon)+\log(1/|a_{1}|^{2})}{\log(\lambda_{1}/\lambda_{2})}\right).$
The adaptive doubling schedule adds only geometric overhead, so total oracle calls remain of the same order. This matches classical power-iteration scaling and confirms that FSPA changes objective robustness (magnitude invariance), not gap-limited asymptotics.
For comparison, we also measure convergence of classical power iteration under identical initialization conditions (Fig. 2).
Figure 2: Empirical validation of the gap-dependent oracle complexity of FSPA. The total number of oracle applications required to reach 99.99% fidelity is plotted against the theoretical scaling variable $1/\log(\lambda_{1}/\lambda_{2})$. Linear regression confirms proportional scaling, consistent with the predicted complexity $\mathcal{O}\!\left(\frac{\log(1/\epsilon)}{\log(\lambda_{1}/\lambda_{2})}\right)$. Classical power iteration is shown for comparison.
Figure 2 strengthens the complexity claim in two ways. First, the near-linear trend against $1/\log(\lambda_{1}/\lambda_{2})$ confirms that the empirical oracle count is controlled by spectral-ratio amplification, not by absolute eigenvalue magnitude. Second, FSPA and classical power iteration follow comparable slopes under matched initialization, indicating that FSPA preserves the same gap-law asymptotics while shifting the algorithmic objective toward projection robustness. Small vertical offsets are consistent with constant-factor effects from scheduling and finite-fidelity stopping criteria rather than a change in asymptotic scaling.
The numerical evidence addresses three points: instability of eigenvector-level targets near degeneracy, magnitude-collapse of estimation-first qPCA, and gap-limited but smooth degradation of projection-first FSPA. For the real-data stability experiment we used the Breast Cancer Wisconsin dataset, a widely used benchmark dataset for machine-learning studies of covariance structure[11, 12]. The dataset contains diagnostic features extracted from digitized images of breast mass tissue and is commonly used to evaluate dimensionality-reduction techniques. For the downstream projection diagnostic we used the Digits dataset, which consists of handwritten digit images represented in a low-dimensional feature space. This dataset provides a controlled environment to evaluate the stability of spectral projection methods under near-degenerate covariance spectra. Both datasets are standard benchmarks available through the UCI Machine Learning Repository [11] and are distributed in the scikit-learn machine learning library [12].
Figure 3: Eigenvector instability versus subspace stability on the Breast Cancer Wisconsin dataset[12]. The real-data covariance matrix is constructed from standardized diagnostic features. Small perturbations strongly rotate individual leading eigenvectors, while dominant-subspace fidelity remains stable. This illustrates why subspace-level metrics are the appropriate object of study in near-degenerate regimes.
Figure 4: Uniform spectral rescaling at fixed gap. Lloyd-style qPCA collapses below a resolution threshold, while FSPA remains stable under global eigenvalue downscaling.
Figure 5: Algorithmic regime map versus spectral gap. Phase-estimation-based methods show a sharp threshold behavior tied to finite resolution; FSPA degrades smoothly as the gap shrinks, consistent with gap-limited amplification.
These data support dominant-subspace fidelity as the relevant metric in near-degenerate settings (Fig. 3), isolate magnitude collapse at fixed gap in estimation-first pipelines (Fig. 4), and show that FSPA replaces threshold-type collapse with continuous gap-limited degradation (Fig. 5). Figure 3 is particularly important for task definition: despite visible rotations of individual leading eigenvectors under small perturbations, the dominant-subspace fidelity remains high, indicating that subspace-level targets are the robust object in realistic near-degenerate regimes. Figure 4 then isolates a different axis—global spectral scale at fixed ordering and gap—and shows that estimation-first behavior can fail abruptly once eigenvalues fall below an effective resolution floor, whereas the normalized FSPA update remains essentially unchanged by the same rescaling. Finally, Fig. 5 clarifies that FSPA does not remove spectral-gap dependence; instead, it converts an abrupt threshold landscape into a smooth degradation profile as the gap decreases, which is the expected signature of amplification-driven projection rather than precision-driven estimation.
Ancillary tests are consistent with the same mechanism; results for weak warm-start bias amplification, perturbative degeneracy lifting inside degenerate dominant subspaces, and downstream Digits projection diagnostics are provided in the Supplemental Material.
FSPA is a projection primitive, not an eigenvalue estimator. Relative to Lloyd-style qPCA and related estimation-first approaches, its differentiating property is update-level invariance to global eigenvalue rescaling, while retaining the same gap-limited convergence structure as power/subspace iteration.[1, 6, 8]
This positioning is deliberate. The method is most natural when the scientific target is dominant-subspace recovery or overlap amplification, especially in nearly degenerate spectra where single-eigenvector targets are unstable. The framework is compatible with block-encoding and QSVT-style polynomial transformations,[9, 10] with normalization interpreted through post-selection or amplitude amplification rather than as a standalone unitary.
In conclusion, this work separates spectral projection from spectral estimation in qPCA regimes and positions FSPA as a projection-first alternative. FSPA is invariant under global eigenvalue rescaling, remains gap-limited with power-iteration-type behavior, and converges to dominant invariant subspaces in degenerate settings. Numerical evidence shows that FSPA avoids the magnitude-collapse behavior seen in estimation-first pipelines without claiming improved gap asymptotics, making it a compact and robust primitive for overlap-amplification tasks.

### 23 Mar 2026 v2

@@
 Quantum ML, Principal Component Analysis, Projection
-preprint: AIP/123-QED
 Quantum principal component analysis (qPCA) is one of the foundational ideas in quantum machine learning (QML) because it connects quantum linear-algebra methods to dimensionality reduction, feature extraction, and covariance-structure analysis in high-dimensional data.[1, 2, 3, 4, 5] Its importance is both conceptual and algorithmic: qPCA helped establish the view that quantum states can be treated as computational objects for learning tasks, rather than merely measured as passive outputs.[1]
@@
   • Eigenvalue estimation: approximate $\lambda_{j}$ to prescribed precision.
+For quick reference, Table 1 summarizes the practical distinction between phase-estimation-driven qPCA and the projection-first FSPA viewpoint.
+Table 1: Direct algorithmic comparison of phase-estimation-driven qPCA (QPE-style) and FSPA.
+Criterion QPE-style qPCA FSPA
+Primary goal Estimation-first Projection-first
+Algorithmic target High-resolution eigenvalue/eigenvector identification Dominant invariant-subspace recovery and overlap amplification
+Scale sensitivity Sensitive to absolute eigenvalue magnitude; resolution-limited collapse can occur under uniform downscaling Invariant under uniform spectral rescaling through normalized updates
+Degeneracy handling Leading-eigenvector targets can be unstable in near-degenerate regimes Stable convergence to the dominant invariant subspace; direction selection only after symmetry lifting
+Required primitive Phase estimation using $e^{-i\rho t}$ with sufficient phase resolution Block-encoding/QSVT polynomial filtering with operational normalization (post-selection/heralding/amplitude amplification)
 FSPA targets dominant-subspace recovery and overlap amplification directly rather than eigenvalue estimation. The iteration used in this work is:
@@
 12:end for
+[…]
 […]
@@
 These propositions clarify the roles of scale invariance, spectral ratios, and degeneracy. The main quantitative guarantee is the following.
-Theorem 1 (Oracle Complexity of FSPA)
+Theorem 1 (Block-encoding query complexity)
 Let $\rho$ be a Hermitian operator with spectral decomposition
@@
 $\mathcal{O}\!\left(\frac{\log\!\left(\frac{1}{\epsilon}\right)+\log\!\left(\frac{1}{|a_{1}|^{2}}\right)}{\log(\lambda_{1}/\lambda_{2})}\right)$
-applications of $\rho$.
+queries to a block-encoding (or equivalent access primitive) for $\rho$.
 Proof sketch: We expand the initial state in the eigenbasis,
@@
 Relative to Lloyd-style qPCA and related estimation-first approaches, its differentiating property is update-level invariance to global eigenvalue rescaling, while retaining the same gap-limited convergence structure as power/subspace iteration.[1, 6, 8]
-This positioning is deliberate.
-The method is most natural when the scientific target is dominant-subspace recovery or overlap amplification, especially in nearly degenerate spectra where single-eigenvector targets are unstable.
+This positioning is intentional.
+The method is best suited to settings in which the scientific objective is dominant-subspace recovery or overlap amplification, particularly for nearly degenerate spectra where single-eigenvector targets are unstable.
 The framework is compatible with block-encoding and QSVT-style polynomial transformations,[9, 10] with normalization interpreted through post-selection or amplitude amplification rather than as a standalone unitary.
@@
 Numerical evidence shows that FSPA avoids the magnitude-collapse behavior seen in estimation-first pipelines without claiming improved gap asymptotics, making it a compact and robust primitive for overlap-amplification tasks.
+Supplementary Information
+Supplemental Figures
+In the following we provide short technical context for each supplemental figure, including what is varied, what metric is reported, and how each plot supports the projection-first interpretation of FSPA.
+Figure S1: Warm-start amplification at small gap. Nonzero initial overlap is amplified efficiently; zero bias does not induce artificial symmetry breaking.
+Figure S1 reports fidelity growth under near-degenerate spectra for different warm-start conditions.
+The horizontal axis is the iteration budget (equivalently oracle applications), and the vertical axis is overlap/fidelity with the target dominant subspace or direction.
+Curves with nonzero initial bias show monotone amplification consistent with gap-limited power-style filtering, whereas the unbiased case remains symmetry-preserving.
+This confirms that FSPA amplifies existing informative overlap without introducing spurious direction selection.
+Figure S2: Perturbative degeneracy lifting. A positive semidefinite perturbation inside a degenerate dominant subspace selects a direction that FSPA recovers with high fidelity.
+Figure S2 studies controlled symmetry breaking inside a degenerate leading subspace via a positive semidefinite perturbation.
+The plot tracks recovery fidelity as a function of perturbation strength and/or iteration depth.
+As soon as a nonzero perturbative preference is present, FSPA converges to the selected direction with high fidelity; in the zero-perturbation limit, behavior is consistent with subspace convergence rather than unique-eigenvector recovery.
+This matches the theoretical distinction between degenerate-subspace projection and direction selection after degeneracy lifting.
+Figure S3: Downstream diagnostic on Digits after top-$k$ projection. Accuracy reflects projection stability; it is not a claim of learning advantage.
+Figure S3 provides an application-level diagnostic using a Digits dataset embedding after top-$k$ spectral projection.
+The reported accuracy is used only as a proxy for projection quality under parameter variation (e.g., global spectral scaling $\alpha$), not as a benchmark claim against dedicated learning pipelines.
+The key observation is stability of downstream behavior in regimes where projection quality is preserved, supporting the practical relevance of magnitude-robust spectral filtering.